Solve the initial value problem





y"-3y'-10y=0, y'(0)=1, y'(0)=2

1
Expert's answer
2022-02-07T13:54:05-0500

"y"-3y'-10y=0, y(0)=1, y'(0)=2"

Auxiliary equation is

"k^2-3k-10=0"

"k=\frac{3\pm \sqrt{9+40}}{2}"

"k_1=-2,k_2=5"

General solution is

"y=c_1e^{-2x}+c_2e^{5x}...(1)"

Since, "y(0)=1"

"1=c_1+c_2...(2)"

And

"y'=-2c_1e^{-2x}+5c_2e^{5x}"

Also "y'(0)=2"

"2=-2c_1+5c_2...(3)"

Multiply equation (2) with 2 and add with equation (3), we get

"4=7c_2\Rightarrow c_2=\frac{4}{7}"

From equation (2),

"1=c_1+\frac 4 7\Rightarrow c_1=\frac 3 7"


"y=\frac 3 7e^{-2x}+\frac 4 7e^{5x}\Rightarrow 7y=3e^{-2x}+4e^{5x}"